# Distribution-informed Online Conformal Prediction

### v1

@@ 1 Introduction
Uncertainty quantification is essential for making reliable forecasts, as prediction errors could lead to severe consequences, particularly in high-risk domains including epidemiology and finance. Despite the availability of various methods for uncertainty quantification, such as confidence calibration [guo2017calibration] and Bayesian networks [fortunato2017bayesian], these approaches often fail to provide provable coverage guarantees, which significantly restricts their applicability and reliability.
@@
To address this limitation, Conformal Prediction (CP) [vovk2005algorithmic] stands out as a powerful technique to construct prediction sets with model-agnostic and distribution-free properties. Under the assumption of data exchangeability, it theoretically guarantees that the prediction sets contain the true label with a specified probability. Consequently, CP has shown promising performance in multiple domains, such as large language models [gui2024conformal], robotics [lindemann2023safe], and image classification [angelopoulos2020uncertainty].
@@
Despite the success of CP for exchangeable data, it remains challenging to directly perform CP in scenarios involving distribution shifts and temporal dependence that violate the exchangeability assumption. To resolve this problem, significant efforts have been made to extend CP methods beyond exchangeability [barber2023conformal, xu2021conformal, tibshirani2019conformal, yang2024doubly].
@@
Recently, a large amount of research is emerging in Conformal Prediction for time series. Pioneered by ACI [aci_gibbs2021adaptive], adversarial online learning methods were incorporated into the CP framework to adapt to arbitrary online distribution shifts. To improve it, Conformal PID [pid_angelopoulos2024conformal] incorporated PID control and ECI [wu2025error] leveraged error quantification via smoothed quantile loss. To further determine the step size for dynamic environments, other online learning techniques were used such as directly choosing decaying step size [angelopoulos2024online], expert aggregation [zaffran2022adaptive, gibbs2024conformal], and strongly adaptive regret [bhatnagar2023improved, hajihashemi2024multi]. However, they focus on fully adversarial environments and adjust prediction sets simply based on the iterative steps, without considering the underlying data patterns. Consequently, despite their well-established theoretical basis, they tend to construct overly conservative prediction sets in practice.
@@
Instead of online learning methods, another line of work [xu2021conformal, xu2023sequential, lee2024kernel] proposed methods to directly obtain prediction sets via estimated distribution function of non-conformity scores, such as SPCI [xu2023sequential], which employed random forest as an estimator. Their theoretical results require assumptions like model consistency and distributional smoothness, and hence are not distribution-free. Other methods focus on assigning weights to historical scores. HOP-CPT [auer2023conformal] utilized the Modern Hopfield Network and CT-SSF [chen2024conformalized] leveraged the inductive bias in deep representation learning to dynamically adjust weights. However, these approaches often lack theoretical results, such as finite-sample coverage, making them potentially unreliable and suffer from overfitting issues.
@@
In this work, we propose Conformal Optimistic Prediction (COP), an online CP algorithm that incorporates underlying data pattern of the non-conformity scores into the update rule. The key step is a refinement update based on the linear approximation of the expected quantile loss and an estimated cumulative distribution function (CDF). COP maintains the well-grounded coverage guarantees inherent to traditional methods, while taking advantage of the distribution information to construct tighter prediction intervals. We also establish a connection to the online optimistic gradient descent [rakhlin2013online, zhao2024adaptivity], with a scaled distribution-informed hint. From this standpoint, we derive a joint bound on regret and coverage, which further confirms the validity of our approach. Moreover, COP does not need to compute inverse of the estimated CDF as [lee2024kernel] and [wang2025mirror] do, thereby reducing computational cost and unnecessary numerical error. Our contributions can be summarized as follows:
@@ 2.1 Problem Setup
However, in an adversarial setting where exchangeability does not hold, such as time series data with strong correlations, it is very difficult to achieve such a real-time coverage guarantee. Recent works [bhatnagar2023improved, pid_angelopoulos2024conformal, angelopoulos2024online, podkopaevadaptive] have turned to consider the following long-term coverage guarantee:
@@ 2.2 Conformal Prediction for Sequential Data
The idea to set $q_{t}$ through online gradient descent with a fixed step size was first introduced in ACI [aci_gibbs2021adaptive]. Subsequent improvements include improving update rules by introducing decaying stepsizes [angelopoulos2024online], PID control [pid_angelopoulos2024conformal], and error quantification [wu2025error]. Moreover, some methods borrow online learning techniques and consider not only coverage guarantee, but also dynamic regret analysis [gibbs2024conformal, bhatnagar2023improved, podkopaevadaptive]. However, these methods only targeted at fully adversarial environments and ignored the predictive part in the sequence. Another line of work aims at learning the distribution of data with estimated distribution functions [xu2023sequential, lee2024kernel] or capturing temporal structure through deep neural networks [chen2024conformalized, auer2023conformal]. However, both of them cannot provide provable coverage guarantees.
Most relate to us are the LQT algorithm proposed by [arecesonline] and the Conformal PID algorithm in [angelopoulos2023prediction]. They consider both adversarial and predictable data. However, the application of LQT relies heavily on linear autoregressive structures. For Conformal PID, the selection of its scorecaster model is arbitrary and lacks principled guidance. An inaccurate or excessively complex scorecaster may impair performance by introducing additional variance.
@@ 2.3 Optimistic Online Gradient Descent
Optimistic online gradient descent (OOGD) has been extensively studied and applied across various domains. It aims at adding prior knowledge about the sequence within the paradigm of online learning. [rakhlin2013online] proposed the general version of the Optimistic Mirror Descent algorithm, which was bulit upon the results in [chiang2012online]. Recently, OOGD has been adopted in Riemannian online convex optimization [wang2023riemannian, roux2025implicit], generative models [daskalakis2017training], reinforcement learning [ito2021parameter, bubeck2019improved], etc. Other theoretical results of OOGD include problem-dependent regret bounds [chen2024optimistic, zhao2024adaptivity], saddle point problem [daskalakis2018limit, mertikopoulos2018optimistic, mokhtari2020unified] and other last iterate convergence properties [gorbunov2022last].
@@ 3.1 Conformal Optimistic Prediction
One critism of eq.˜3 is that the good coverage is obtained due to reactively correcting past mistakes and through a cancellation of positive and negative errors. The radius $q_{t}$ may exhibit significant fluctuations around the true value [aci_gibbs2021adaptive, wu2025error], leading to inefficient prediction intervals. An alternative way is to directly estimate the conditional quantiles or distribution function $\hat{F}_{t}$ of the scores in an online fashion [lee2024kernel]. Then $q_{t}$ can directly take $\hat{F}^{-1}_{t}(1-\alpha)$. However, the validity of these methods fully counts on the estimated quantiles or CDF, which may suffer from overfitting when addressing adversarial data.
@@
We remark that our refinement can also be implemented on another widely employed online CP method, adaptive conformal inference (ACI) [aci_gibbs2021adaptive], as detailed in Section˜A.2. There are many choices for $\hat{F}_{t+1}$, such as empirical cumulative distribution function (ECDF) and kernel-based estimation [silverman1986density]. As a brief clarification on the benefit of our refinement, we present the proposition as follows.
@@ 3.2 Connection with OOGD
In the literature of OOGD, $M_{t+1}$ is the optimistic term, aiming at guessing the next move and incorporating it into the objective [rakhlin2013online]. While classical OOGD often takes $M_{t+1}$ as the previous gradient, COP uses a distribution-informed hint:
@@
Besides the coverage guarantee in Section˜3.3, dynamic regret is often considered as an additional performance metric for online CP [bhatnagar2023improved, ramalingamrelationship]. From the view of OOGD, we obtain a joint regret-coverage bound with constant learning rate in Theorem˜1. The result for dynamic learning rates can be seen in Section˜B.2.
@@
The term $\sum_{t=1}^{T}\frac{1}{2\eta}\left(\|u_{t}-\hat{q}_{t}\|_{2}^{2}-\|u_{t-1}-\hat{q}_{t}\|_{2}^{2}\right)$ reflects the non-stationarity of environments [zhao2024adaptivity]. Although without distributional assumption, regret bounds and coverage bounds of online CP do not imply each other [angelopoulos2025gradient], Theorem˜1 shows that properly choosing $M_{t}$ can simultaneously reduce the bound for both regret and coverage. Specifically, note that $\mathbb{E}\|\alpha-err_{t}-M_{t}\|_{2}^{2}|\mathcal{S}_{t-1}\approx\mathbb{E}\|\alpha-\mathds{1}(s_{t}>\hat{q}_{t})-M_{t}\|_{2}^{2}|\mathcal{S}_{t-1}$, a natural idea for sequentially choosing $M_{t}$ would be taking it close to:
@@ 3.3 Coverage Guarantees
Note that, $\|\Delta_{1:T}\|_{1}$ is closely related to the number of times we increase the step size. Specifically, $\|\Delta_{1:T}\|_{1}\leq 2N_{t}/(\min_{t\leq T}\eta_{t})$, where $N_{T}=\sum_{t=1}^{T}\mathds{1}(\eta_{t+1}>\eta_{t})$. Hence, the upper bound in Theorem˜2 will converge to zero as long as $N_{T}/(\min_{t\leq T}\eta_{t})=o(T)$. In particular, if we consider decaying $\eta_{t}$ in [angelopoulos2024online], then $\Omega_{T}=\eta_{1},\|\Delta_{1:T}\|_{1}=1/\eta_{T},\|\Delta_{1:T}\|_{1}/T\leq N_{T}/T\eta_{T}$. Setting $\eta_{T}=o(N_{T}/T)$ ensures the long-term coverage in eq.˜1.
@@
Choosing constant learning rates in eq.˜3 will cause oscillations [gibbs2024conformal]. Theorem˜3 demonstrates that, under the i.i.d. assmptions and choosing learning rates satisfying the Robbins-Monro condition [robbins1951stochastic], $q_{t}$ exhibits a convergence behavior, even if we add bounded optimistic terms $M_{t}$ in every update.
@@ 4.1 Setup
Datasets. We evaluate five simulation datasets under changepoints, distribution drift [barber2023conformal], variance changepoint, heavy-tailed noise, and extreme distribution drift. Besides, we evaluate four real-world datasets: Amazon stock, Google stock [cam2018stock], electricity demand [harries1999splice] and temperature in Delhi [sumanth2017climate]. In the subsequent sections, we will provide a detailed introduction to each of these datasets.
@@
Baselines. We compare with seven state-of-the-art methods: ACI [aci_gibbs2021adaptive], OGD, SF-OGD [bhatnagar2023improved], decay-OGD [angelopoulos2024online], Conformal PID [pid_angelopoulos2024conformal], ECI [wu2025error] and LQT(fixed) [arecesonline]. Several follow-up works on ACI consider adaptively setting learning rates via the expert aggregation technique [gibbs2024conformal, bhatnagar2023improved]. COP and the seven baselines are orthogonal to these works and can be naturally incorporated by serving as a single expert.
@@
Hyperparameters. The proposed COP has three hyperparameters, the base learning rate $\eta$, scale factor $\lambda=0.5$ and the window length $w=100$. Same as previous works, the adaptive learning rates $\eta_{t}=\eta\cdot(\max\{s_{t-w+1},\dots,s_{t}\}-\min\{s_{t-w+1},\dots,s_{t}\})$. For reproducibility, all baseline implementations leverage the open-source Python codes from [wu2025error] or [arecesonline]. The hyperparameters of LQT need to be tuned via grid search. The operational ranges of $\eta$ and more details about $\eta_{t}$ for each methods can be found in Appendix˜D.
@@ 4.2 Simulation Dataset
Distribution
Drift
ACI 89.9 $\infty$ 6.69 89.8 $\infty$ 6.56 89.9 $\infty$ 6.79
@@ 4.3 Real-world Datasets
Amazon
Stock
ACI 90.2 $\infty$ 46.97 89.8 $\infty$ 13.77 89.7 $\infty$ 12.31
@@
Google
Stock
ACI 90.0 $\infty$ 66.83 89.8 $\infty$ 18.64 90.5 $\infty$ 32.78
@@
Electricity
Demand
ACI 90.1 $\infty$ 0.443 90.1 $\infty$ 0.105 90.2 $\infty$ 0.055

### v2

@@ 1 Introduction
Uncertainty quantification is essential for making reliable forecasts, as prediction errors could lead to severe consequences, particularly in high-risk domains including epidemiology and finance. Despite the availability of various methods for uncertainty quantification, such as confidence calibration (Guo et al., 2017) and Bayesian networks (Fortunato et al., 2017), these approaches often fail to provide provable coverage guarantees, which significantly restricts their applicability and reliability.
@@
To address this limitation, Conformal Prediction (CP) (Vovk et al., 2005) stands out as a powerful technique to construct prediction sets with model-agnostic and distribution-free properties. Under the assumption of data exchangeability, it theoretically guarantees that the prediction sets contain the true label with a specified probability. Consequently, CP has shown promising performance in multiple domains, such as large language models (Gui et al., 2024), robotics (Lindemann et al., 2023), image classification (Angelopoulos et al., 2020), and decision-making (Bao et al., 2025).
@@
Despite the success of CP for exchangeable data, it remains challenging to directly perform CP in scenarios involving distribution shifts and temporal dependence that violate the exchangeability assumption. To resolve this problem, significant efforts have been made to extend CP methods beyond exchangeability (Barber et al., 2023; Xu and Xie, 2021; Tibshirani et al., 2019; Yang et al., 2024).
@@
Recently, a large amount of research is emerging in Conformal Prediction for time series. Pioneered by ACI (Gibbs and Candès, 2021), adversarial online learning methods were incorporated into the CP framework to adapt to arbitrary online distribution shifts. To improve it, Conformal PID (Angelopoulos et al., 2023b) incorporated PID control and ECI (Wu et al., 2025) leveraged error quantification via smoothed quantile loss. To further determine the step size for dynamic environments, other online learning techniques were used such as directly choosing decaying step size (Angelopoulos et al., 2024), expert aggregation (Zaffran et al., 2022; Gibbs and Candès, 2024), and strongly adaptive regret (Bhatnagar et al., 2023; Hajihashemi and Shen, 2024). However, they focus on fully adversarial environments and adjust prediction sets simply based on the iterative steps, without considering the underlying data patterns. Consequently, despite their well-established theoretical basis, they tend to construct overly conservative prediction sets in practice.
@@
Instead of online learning methods, another line of work (Xu and Xie, 2021; 2023; Lee et al., 2024) proposed methods to directly obtain prediction sets via estimated distribution function of non-conformity scores, such as SPCI (Xu and Xie, 2023), which employed random forest as an estimator. Their theoretical results require assumptions like model consistency and distributional smoothness, and hence are not distribution-free. Other methods focus on assigning weights to historical scores. HOP-CPT (Auer et al., 2023) utilized the Modern Hopfield Network and CT-SSF (Chen et al., 2024a) leveraged the inductive bias in deep representation learning to dynamically adjust weights. However, these approaches often lack theoretical results, such as finite-sample coverage, making them potentially unreliable and suffer from overfitting issues.
@@
In this work, we propose Conformal Optimistic Prediction (COP), an online CP algorithm that incorporates underlying data pattern of the non-conformity scores into the update rule. The key step is a refinement update based on the linear approximation of the expected quantile loss and an estimated cumulative distribution function (CDF). COP maintains the well-grounded coverage guarantees inherent to traditional methods, while taking advantage of the distribution information to construct tighter prediction intervals. We also establish a connection to the online optimistic gradient descent (Rakhlin and Sridharan, 2013; Zhao et al., 2024), with a scaled distribution-informed hint. From this standpoint, we derive a joint bound on regret and coverage, which further confirms the validity of our approach. Moreover, COP does not need to compute inverse of the estimated CDF as Lee et al. (2024) and Wang et al. (2025) do, thereby reducing computational cost and unnecessary numerical error. Our contributions can be summarized as follows:
@@ 2.1 Problem Setup
However, in an adversarial setting where exchangeability does not hold, such as time series data with strong correlations, it is very difficult to achieve such a real-time coverage guarantee. Recent works (Bhatnagar et al., 2023; Angelopoulos et al., 2023b; 2024; Podkopaev et al., 2024) have turned to consider the following long-term coverage guarantee:
@@ 2.2 Conformal Prediction for Sequential Data
The idea to set $q_{t}$ through online gradient descent with a fixed step size was first introduced in ACI (Gibbs and Candès, 2021). Subsequent improvements include improving update rules by introducing decaying stepsizes (Angelopoulos et al., 2024), PID control (Angelopoulos et al., 2023b), and error quantification (Wu et al., 2025). Moreover, some methods borrow online learning techniques and consider not only coverage guarantee, but also dynamic regret analysis (Gibbs and Candès, 2024; Bhatnagar et al., 2023; Podkopaev et al., 2024). Bao et al. (2024) investigated online selective conformal inference problem for i.i.d. data stream. However, these methods only targeted at fully adversarial environments and ignored the predictive part in the sequence. Another line of work aims at learning the distribution of data with estimated distribution functions (Xu and Xie, 2023; Lee et al., 2024) or capturing temporal structure through deep neural networks (Chen et al., 2024a; Auer et al., 2023). However, both of them cannot provide provable coverage guarantees.
Most relate to us are the LQT algorithm proposed by Areces et al. (2025) and the Conformal PID algorithm in Angelopoulos et al. (2023a). They consider both adversarial and predictable data. However, the application of LQT relies heavily on linear autoregressive structures. For Conformal PID, the selection of its scorecaster model is arbitrary and lacks principled guidance. An inaccurate or excessively complex scorecaster may impair performance by introducing additional variance.
@@ 2.3 Optimistic Online Gradient Descent
Optimistic online gradient descent (OOGD) has been extensively studied and applied across various domains. It aims at adding prior knowledge about the sequence within the paradigm of online learning. Rakhlin and Sridharan (2013) proposed the general version of the Optimistic Mirror Descent algorithm, which was bulit upon the results in (Chiang et al., 2012). Recently, OOGD has been adopted in Riemannian online convex optimization (Wang et al., 2023; Roux et al., 2025), generative models (Daskalakis et al., 2017), reinforcement learning (Ito, 2021; Bubeck et al., 2019), etc. Other theoretical results of OOGD include problem-dependent regret bounds (Chen et al., 2024b; Zhao et al., 2024), saddle point problem (Daskalakis and Panageas, 2018; Mertikopoulos et al., 2018; Mokhtari et al., 2020) and other last iterate convergence properties (Gorbunov et al., 2022).
@@ 3.1 Conformal Optimistic Prediction
One critism of eq.˜3 is that the good coverage is obtained due to reactively correcting past mistakes and through a cancellation of positive and negative errors. The radius $q_{t}$ may exhibit significant fluctuations around the true value (Gibbs and Candès, 2021; Wu et al., 2025), leading to inefficient prediction intervals. An alternative way is to directly estimate the conditional quantiles or distribution function $\hat{F}_{t}$ of the scores in an online fashion (Lee et al., 2024). Then $q_{t}$ can directly take $\hat{F}^{-1}_{t}(1-\alpha)$. However, the validity of these methods fully counts on the estimated quantiles or CDF, which may suffer from overfitting when addressing adversarial data.
@@
We remark that our refinement can also be implemented on another widely employed online CP method, adaptive conformal inference (ACI) (Gibbs and Candès, 2021), as detailed in Section˜A.2. There are many choices for $\hat{F}_{t+1}$, such as empirical cumulative distribution function (ECDF) and kernel-based estimation (Silverman, 1986). As a brief clarification on the benefit of our refinement, we present the proposition as follows.
@@ 3.2 Connection with OOGD
In the literature of OOGD, $M_{t+1}$ is the optimistic term, aiming at guessing the next move and incorporating it into the objective (Rakhlin and Sridharan, 2013). While classical OOGD often takes $M_{t+1}$ as the previous gradient, COP uses a distribution-informed hint:
@@
Besides the coverage guarantee in Section˜3.3, dynamic regret is often considered as an additional performance metric for online CP (Bhatnagar et al., 2023; Ramalingam et al., 2025). From the view of OOGD, we obtain a joint regret-coverage bound with constant learning rate in Theorem˜1. The result for dynamic learning rates can be seen in Section˜B.2.
@@
The term $\sum_{t=1}^{T}\frac{1}{2\eta}\left(\|u_{t}-\hat{q}_{t}\|_{2}^{2}-\|u_{t-1}-\hat{q}_{t}\|_{2}^{2}\right)$ reflects the non-stationarity of environments (Zhao et al., 2024). Although without distributional assumption, regret bounds and coverage bounds of online CP do not imply each other (Angelopoulos et al., 2025), Theorem˜1 shows that properly choosing $M_{t}$ can simultaneously reduce the bound for both regret and coverage. Specifically, note that $\mathbb{E}\|\alpha-err_{t}-M_{t}\|_{2}^{2}|\mathcal{S}_{t-1}\approx\mathbb{E}\|\alpha-\mathds{1}(s_{t}>\hat{q}_{t})-M_{t}\|_{2}^{2}|\mathcal{S}_{t-1}$, a natural idea for sequentially choosing $M_{t}$ would be taking it close to:
@@ 3.3 Coverage Guarantees
Note that, $\|\Delta_{1:T}\|_{1}$ is closely related to the number of times we increase the step size. Specifically, $\|\Delta_{1:T}\|_{1}\leq 2N_{t}/(\min_{t\leq T}\eta_{t})$, where $N_{T}=\sum_{t=1}^{T}\mathds{1}(\eta_{t+1}>\eta_{t})$. Hence, the upper bound in Theorem˜2 will converge to zero as long as $N_{T}/(\min_{t\leq T}\eta_{t})=o(T)$. In particular, if we consider decaying $\eta_{t}$ in (Angelopoulos et al., 2024), then $\Omega_{T}=\eta_{1},\|\Delta_{1:T}\|_{1}=1/\eta_{T},\|\Delta_{1:T}\|_{1}/T\leq N_{T}/T\eta_{T}$. Setting $\eta_{T}=o(N_{T}/T)$ ensures the long-term coverage in eq.˜1.
@@
Choosing constant learning rates in eq.˜3 will cause oscillations (Gibbs and Candès, 2024). Theorem˜3 demonstrates that, under the i.i.d. assmptions and choosing learning rates satisfying the Robbins-Monro condition (Robbins and Monro, 1951), $q_{t}$ exhibits a convergence behavior, even if we add bounded optimistic terms $M_{t}$ in every update.
@@ 4.1 Setup
Datasets. We evaluate five simulation datasets under changepoints, distribution drift (Barber et al., 2023), variance changepoint, heavy-tailed noise, and extreme distribution drift. Besides, we evaluate four real-world datasets: Amazon stock, Google stock (Nguyen, 2018), electricity demand (Harries et al., 1999) and temperature in Delhi (Vrao., 2017). In the subsequent sections, we will provide a detailed introduction to each of these datasets.
@@
Baselines. We compare with seven state-of-the-art methods: ACI (Gibbs and Candès, 2021), OGD, SF-OGD (Bhatnagar et al., 2023), decay-OGD Angelopoulos et al. (2024), Conformal PID (Angelopoulos et al., 2023b), ECI (Wu et al., 2025) and LQT(fixed) (Areces et al., 2025). Several follow-up works on ACI consider adaptively setting learning rates via the expert aggregation technique (Gibbs and Candès, 2024; Bhatnagar et al., 2023). COP and the seven baselines are orthogonal to these works and can be naturally incorporated by serving as a single expert.
@@
Hyperparameters. The proposed COP has three hyperparameters, the base learning rate $\eta$, scale factor $\lambda=0.5$ and the window length $w=100$. Same as previous works, the adaptive learning rates $\eta_{t}=\eta\cdot(\max\{s_{t-w+1},\dots,s_{t}\}-\min\{s_{t-w+1},\dots,s_{t}\})$. For reproducibility, all baseline implementations leverage the open-source Python codes from Wu et al. (2025) or Areces et al. (2025). The hyperparameters of LQT need to be tuned via grid search. The operational ranges of $\eta$ and more details about $\eta_{t}$ for each methods can be found in Appendix˜D.
@@ 4.2 Simulation Dataset
Distribution Drift ACI 89.9 $\infty$ 6.69 89.8 $\infty$ 6.56 89.9 $\infty$ 6.79
@@ 4.3 Real-world Datasets
Amazon Stock ACI 90.2 $\infty$ 46.97 89.8 $\infty$ 13.77 89.7 $\infty$ 12.31
@@
Google Stock ACI 90.0 $\infty$ 66.83 89.8 $\infty$ 18.64 90.5 $\infty$ 32.78
@@
Electricity Demand ACI 90.1 $\infty$ 0.443 90.1 $\infty$ 0.105 90.2 $\infty$ 0.055
@@ 5 Conclusions
ACKNOWLEDGEMENT
We would like to thank all anonymous reviewers and the area chair for their helpful comments. Shu-Tao Xia was supported in part by the National Natural Science Foundation of China under Grant 62571298. Changliang Zou was supported by the National Key R&D Program of China (Grant Nos. 2022YFA1003703, 2022YFA1003800), and the National Natural Science Foundation of China (Grant No. 12231011).